PUT INTO PYTHON CODE

#include <stdio.h>
#include <float.h>

int main(void) {
float a = 1;
float b = 0.5;
float c = a + b;
int bits = 0;
while (c != a) {
bits = bits + 1;
b = b / 2;
c = a + b;
}
printf("%d\n",bits);
return 0;
}
Answer #1

a=1.0 b=0.5 C=a+b bits=0 while(c!=a): bits= bits+1 b= (b/2) C= a+b print(bits)

a=1.0

b=0.5

c=a+b

bits=0

while(c!=a):

  bits= bits+1

  b= (b/2)

  c= a+b

print(bits)

This is the code

Please comment if you have any query and please like